# Improving Signed Propagation of Graph Neural Network Under Multiple Classes

### v6

@@
Abstract
Graph Neural Networks (GNNs) exhibit satisfactory performance on homophilic networks, where most edges connect two nodes with the same label. However, their effectiveness diminishes as the graphs become heterophilic (low homophily), prompting the exploration of various message-passing schemes. In particular, assigning negative weights to heterophilic edges (signed propagation) for message-passing has gained significant attention, and some studies theoretically confirm its effectiveness. Nevertheless, prior theorems assume binary classification scenarios, which may not hold well for graphs with multiple classes. To solve this limitation, we offer new theoretical insights into GNNs in multi-class environments and identify the drawbacks of employing signed propagation from two perspectives: message-passing and parameter update. We found that signed propagation without considering feature distribution can degrade the separability of dissimilar neighbors, which also increases prediction uncertainty (e.g., conflicting evidence) that can cause instability. To address these limitations, we introduce two novel calibration strategies aiming to improve discrimination power while reducing entropy in predictions. Through theoretical and extensive experimental analysis, we demonstrate that the proposed schemes enhance the performance of both signed and general message-passing neural networks.
Index Terms:
@@ Abstract
I Introduction
@@
The recent proliferation of graph-structured datasets has ignited rapid advancements in graph mining techniques. Especially, Graph Neural Networks (GNNs) provide satisfactory performances in various downstream tasks including node classification and link prediction. The main component of GNNs is message-passing [1], where information is propagated between adjacent nodes and aggregated. Additionally, the entailment of a structural property enhances the representation and the discrimination powers of GNNs [2, 3].
@@
Early GNNs assume the network homophily, where nodes of similar features make connections based on social influence and/or selection theories. Spectral GNNs [1, 2] employ Laplacian smoothing called low-pass filtering to receive low-frequency signals from neighbor nodes. Consequently, these methods fail to adequately address heterophilous graphs [4] and show dismal performance [5] in several heterophilous cases. To resolve this problem, numerous clever algorithms (spatial GNNs) have been proposed, including adjustment of edge coefficients [3, 6] and aggregation of remote but highly similar nodes [7, 8]. However, most of these methods assume positive or unsigned edges, which may impoverish the separation power [9].
Recently, several studies [10, 11] have applied message passing over negative (signed propagation) edges to preserve high-frequency signal exchanges. From the viewpoint of gradient flow, [12] shows that negative eigenvalues can intensify high-frequency signals during propagation. On the other hand, [13] prevents message diffusion by assigning zero-weights (blocking information) to heterophilic edges to alleviate local smoothing. Based on this observation, we raise a question: does signed propagation always attain the best performance on heterophilic graphs?
@@
To address the above question, we focus on the recent investigations [5, 9] that provide valuable insights on message-passing. These studies mathematically analyze the effect of unsigned and signed messages on separability by comparing node feature distributions before and after message passing and aggregation. Furthermore, they reveal that signed messaging always enhances performance on heterophilic graphs. However, our empirical analysis (Fig. 3) shows that this assertion may not hold well under real-world graphs. This discrepancy arises because the prior work assumes binary class graphs [14], impeding their extension to multi-class graphs.
In this paper, we deal with signed messaging on multi-class graphs. We first demonstrate that blind application of signed messaging to multi-class graphs may incur significant performance degradation. To overcome the problem, we propose to utilize two types of calibration [15, 16], which are simple yet effective in enhancing the performance of signed GNNs. In summary, our contributions can be described as follows:
@@
  • We hypothesize and prove that signed propagation may reduce the discrimination power in certain cases while increasing prediction uncertainties. Expanding this understanding, we propose novel strategies based on two types of calibration that are shown to be effective for both signed and unsigned GNNs.
@@
  • We perform extensive experiments with real-world benchmark datasets to validate our theorems. Our experiments demonstrate the effectiveness of the proposed calibration techniques.
II Related Work
@@
Graph Neural Networks (GNNs). Aggregating the information from adjacent nodes, GNNs have shown great potential under semi-supervised learning settings. Early study [1] focused on the spectral graph analysis (e.g., Laplacian decomposition) in a Fourier domain. However, it suffers from large computational costs as the graph scale increases. GCN [2] reduced the overhead by harnessing the localized spectral convolution through the first-order approximation of a Chebyshev polynomial. Another notable approach is spatial-based GNNs [3, 6, 17, 18] which aggregate information in a Euclidean domain. Early spatial techniques have led to the development of many powerful schemes that encompass remote nodes as neighbors.
@@
GNNs on Heterophilic Graphs. Traditional message-passing GNNs fail to perform well in heterophilic graphs [7]. To redeem this problem, recent studies have focused on handling disassortative edges [19, 20, 21, 22] by capturing node differences or by incorporating similar remote nodes as neighbors. More recently, H2GCN [4] suggests the separation of ego and neighbors during aggregation. As another branch, selecting neighbors from non-adjacent nodes [8], configuring path-level pattern [23], finding a compatibility matrix [24], employing adaptive propagation [25], and choosing appropriate architectures [26] have been recently proposed. Some methodologies change the sign of disassortative edges from positive to negative [10, 11, 27, 28] while others assign zero-weights to disassortative edges [13]. Even though [29] demonstrates that signed propagation can effectively handle binary class graphs, further investigation may be needed before applying the same techniques to multi-class graphs.
III Preliminary
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$ be a graph with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges. The node attribute matrix is $X\in\mathbb{R}^{n\times F}$, where $F$ is the dimension of an input vector. Given $X$, the hidden representation of node features $H^{(l)}$ at the l-th layer is derived through message-passing. Here, node $i^{\prime}s$ feature is defined as $h_{i}^{(l)}$. The structural property of $\mathcal{G}$ is represented by its adjacency matrix $A\in\{0,1\}^{n\times n}$. A diagonal matrix $D$ of node degrees is derived from $A$ as $d_{ii}=\sum^{n}_{j=1}{A_{ij}}$. Each node has its label $Y\in\mathbb{R}^{n\times C}$ ($C$ represents the number of classes). Lastly, the global edge homophily ratio, $\mathcal{H}_{g}$, is defined as:
$\mathcal{H}_{g}\equiv\frac{\sum_{(i,j)\in\mathcal{E}1(Y_{i}=Y_{j})}}{|\mathcal{E}|}$ (1)
Likewise, the local homophily ratio, $b_{i}$, of node $i$ is given as:
$b_{i}\equiv\frac{{\sum^{n}_{j=1}A_{ij}\cdot 1(Y_{i}=Y_{j})}}{d_{ii}}$ (2)
@@ IV Theoretical Analysis
We first discuss the mechanism of GNNs in terms of message-passing and parameter update (§IV-A). Then, we introduce previous analyses of signed message passing on binary class graphs (§IV-B) and highlight some misunderstandings revealed through an empirical study with real-world datasets (§IV-C). Based on this observation, we extend the prior analysis to a multi-class scenario and point out some drawbacks of signed messaging in more complicated environments (§IV-D).
IV-A Message-Passing Neural Networks
@@
Generally, GNNs employ alternate steps of propagation and aggregation recursively, during which the node features are updated iteratively. This is widely known as message-passing, which can be represented as follows:
$\begin{gathered}H^{(l+1)}=\phi(\bar{H}^{(l+1)}),\,\,\,\bar{H}^{(l+1)}=AH^{(l)}W^{(l)}\end{gathered}$ (3)
@@
Here, $H^{(0)}=X$ is the initial vector and $H^{(l)}$ is nodes’ hidden representations at the l-th layer. $\bar{H}^{(l+1)}$ is retrieved through message-passing ($A$) and we obtain $H^{(l+1)}$ after an activation function $\phi$. $W^{(l)}$ is trainable weight matrices shared by all nodes. The final prediction is produced by applying cross-entropy $\sigma(\cdot)$ (e.g., log-softmax) to $\bar{H}^{(L)}$ and the loss function is defined as below:
$\mathcal{L}_{GNN}=\mathcal{L}_{nll}(Y,\widehat{Y}),\,\,\widehat{Y}=\sigma(\bar{H}^{(L)})$ (4)
Figure 1: We take an example to illustrate the distribution of node features under (a) binary and (b) multi-class scenarios. Figure (c) represents the aggregation of neighboring nodes ($k_{1},k_{2}$) under multiple classes, where the scale of aggregated vector $k^{\prime}$ is always smaller than $|\mu|$
@@
The parameters are updated by computing negative log-likelihood loss $\mathcal{L}_{nll}$ between the predictions ($\widehat{Y}$) and true labels ($Y$). Since most GNNs assume homophily, they tend to preserve the low-frequency information (local smoothing) [30]. As a result, they fail to pinpoint the differences between heterophilic neighbors and show dismal performance [31, 7]. Recent studies [10, 11, 9, 32] focus on the extraction of high-frequency signals by flipping the sign of disassortative edges from positive to negative. In the following sections, we analyze the influence of information propagation from the perspectives of message-passing (Eq. 3) and parameter update (Eq. 4).
IV-B Using Signed Messages on Binary Classes
Before delving into the multi-class graphs, we assess the implications of signed propagation with graphs containing only two types of classes [29, 9].
@@ IV-B Using Signed Messages on Binary Classes
Message-passing (Signed propagation always has a better discrimination power than all positive propagation) For brevity, we take GCN [2] as an example to inspect how message passing alters node feature vectors. Assuming a binary label ($y_{i}\in$0, 1}), let us inherit several assumptions and notations from [9] as follows: (1) For all nodes $i=\{1,...,n\}$, their degrees $\{d_{i}\}$ and features $\{h_{i}\}$ are i.i.d. random variables. (2) Every class has the same population. (3) With a slight abuse of notation, assume that $h^{(0)}=XW^{(0)}$ is the first layer projection of the initial node features. (4) The node feature of label $y_{i}$ follows the distribution ($\mu$ or $-\mu$) as,
$\begin{gathered}\mathbb{E}(h^{(0)}_{i}|y_{i})=\begin{cases}\mu,&\text{if }y_{i}=0\\ -\mu,&\text{if }y_{i}=1\end{cases}\end{gathered}$ (5)
@@
We first introduce the equations from prior work [9] (Eq. IV-B and Eq. IV-B), which demonstrate the updates in the expectations of node $i$’s latent features $\mathbb{E}(h^{(0)}_{i}|y_{i})$ after a single-hop application of message passing $\mathbb{E}(h^{(1)}_{i}|y_{i},d_{i})$ below:
$h^{(1)}_{i}={h^{(0)}_{i}\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}{h^{(0)}_{j}\over\sqrt{(d_{i}+1)(d_{j}+1)}}$ (6)
@@
(Vanilla GCN) As illustrated in Figure 1a (binary class), we assume $h_{i}\sim N(\mu,0)$ for $y_{i}=0$. Otherwise, $h_{i}\sim N(-\mu,0)$. Let us assume that $1/d_{ij}=1/\sqrt{(d_{i}+1)(d_{j}+1)}$. Then, the updated node feature $\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})$ derived through the message-passing of vanilla GCN (Eq. 6) is as follows:
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={\mu\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}\left({b_{i}\over d_{ij}}\mu-{(1-b_{i})\over d_{ij}}\mu\right)$
$=\left({1\over d_{i}+1}+{2b_{i}-1\over d_{i}+1}\sum_{j\in\mathcal{N}_{i}}{\sqrt{d_{i}+1}\over\sqrt{d_{j}+1}}\right)\mu$
$=\left({1+(2b_{i}-1)d^{\prime}_{i}\over d_{i}+1}\right)\mu\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,*d^{\prime}_{i}=\sum_{j\in\mathcal{N}_{i}}{1\over d_{ij}}$ (7)
@@
(Signed GCN) To define the mechanism of signed GCN, we need to assume an error ratio $e$ that stands for the accuracy ($1-e$) of configuring the sign of heterophilous edges. Considering that the sign of heterophilous nodes is flipped inaccurately with a probability $e$, the update of the signed GCN can be defined as below:
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={\mu\over d_{i}+1}+$
$\sum_{j\in\mathcal{N}_{i}}\left({\mu(1-e)-\mu e\over d_{ij}}b_{i}+{\mu(1-e)-\mu e\over d_{ij}}(1-b_{i})\right)$
$=\left({1\over d_{i}+1}+{1-2e\over d_{i}+1}\sum_{j\in\mathcal{N}_{i}}{1\over d_{ij}}\right)\mu$
$=\left({1+(1-2e)d^{\prime}_{i}\over d_{i}+1}\right)\mu$ (8)
@@
Theorem 4.1 (Zero-weight GCN).
Given an error ratio $e$, assigning zero weights to heterophilic edges leads to the following update in the expectation of feature distribution:
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={\mu\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}\left({\mu(1-e)\over d_{ij}}b_{i}-{\mu e\over d_{ij}}(1-b_{i})\right)$
$=\left({1+(b_{i}-e)d^{\prime}_{i}\over d_{i}+1}\right)\mu.$ (9)
For all message-passing schemes, if the coefficient of $\mathbb{E}(h^{(0)}_{i}|y_{i})$ is smaller than 1, node feature vectors move towards the decision boundary and message-passing loses its discrimination power [9]. Based on this, we compare the separability of signed GCN and zero-weight GCN below.
@@
Omitting the overlapping part of Eq. IV-B and Eq. 4.1, their difference Z can be derived using the error ratio and the local homophily ratio as,
$\begin{gathered}Z=(1-2e)-(b_{i}-e)=1-e-b_{i},\end{gathered}$ (10)
@@
(a) Binary class case
(b) Multi-class case
@@
Figure 2: We plot the Z in (a) Eq. 10 and (b) Eq. 18 to compare the discrimination powers of signed GCN and zero-weight GCNs. The red and blue colored parts indicate the regions where signed GCN and zero-weight GCN produce superior performances, respectively
Parameter update (The feature vectors of two nodes connected with a negative edge are separated farther during training) Let us assume that an ego ($i$) and one of its neighbor nodes ($j$) are connected with a negative edge. Please ignore other neighbors to concentrate on the separation effect (sign) of propagated messages. The ego node $i$’s feature vector is updated after message-passing as below:
@@
$\begin{gathered}\widehat{Y}_{i}=\sigma(\bar{H}_{i}^{L})=\sigma({\bar{H}_{i}^{(L)}\over d_{i}+1}-{\bar{H}_{j}^{(L)}\over\sqrt{(d_{i}+1)(d_{j}+1)}})\end{gathered}$ (11)
@@
Given that $Y_{i}=k$, the loss ($\mathcal{L}_{nll}$) between the true label ($Y_{i}\in\mathcal{R}^{C}$) and a prediction ($\widehat{Y}_{i}\in\mathcal{R}^{C}$) is defined as below:
$\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})=-\log(\widehat{y}_{i,k})$ (12)
@@
Accordingly, the update procedure is given as $\widehat{y}^{(t+1)}_{i,k}=\widehat{y}^{t}_{i,k}-\eta\nabla_{i}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})$ and $\widehat{y}^{(t+1)}_{j,k}=\widehat{y}^{t}_{j,k}-\eta\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})$. Here, $\eta$ is the learning rate and $\nabla$ represents a partial derivative. The gradient of node $j$, $\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}$, is given by:
@@
$\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}={\partial\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}\over\partial\widehat{y}_{i,k}}={\partial\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}\over\partial\widehat{y}_{i,k}}\cdot{\partial\widehat{y}_{i,k}\over\partial h^{(L)}_{\textbf{j},k}}$
$=-{1\over\widehat{y}_{i,k}}\cdot(\widehat{y}_{i,k}(1-\widehat{y}_{i,k})(-1))=1-\widehat{y}_{i,k}>0$ (13)
@@
Since the column-wise components of the last weight matrix $W^{(L)}$ (Eq. 3) act as independent classifiers, we can infer that as the training epoch ($t$) progresses, the probability of class $k$ for node $j$ ($\widehat{y}_{j,k}$) decreases since $-\eta\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})<0.$

### v7

# Improving Signed Propagation for Graph Neural Networks in Multi-Class Environments

@@
Abstract.
Graph Neural Networks (GNNs) perform well on homophilic networks, where edges generally connect nodes with the same label. However, their performance declines on heterophilic networks, prompting the investigation of alternative message-passing approaches. Notably, assigning negative weights to heterophilic edges (signed propagation) has gained attention, with some studies confirming its effectiveness theoretically. However, these studies often assume binary classification, which may not generalize to multi-class graphs. To address this, we provide new theoretical insights into GNNs in multi-class settings, highlighting the limitations of signed propagation regarding both message-passing and parameter updates. Our findings show that ignoring feature distribution in signed propagation can reduce the separability of dissimilar neighbors, increasing prediction uncertainty and instability. To overcome these challenges, we propose two novel calibration strategies that enhance discrimination and reduce entropy in predictions. Both theoretical and extensive experimental analyses demonstrate that our methods improve the performance of signed and general GNNs.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024conference: under review; ; booktitle: under reviewprice: 15.00isbn: 978-1-4503-9236-5/22/10doi: 10.1145/ccs: Computing methodologies Machine learning
1. Introduction
@@
The recent proliferation of graph-structured datasets has ignited rapid advancements in graph mining techniques. Graph Neural Networks (GNNs) provide satisfactory performances in various downstream tasks including node classification and link prediction. The main component of GNNs is message-passing (Defferrard et al., 2016), where information is propagated between adjacent nodes and aggregated. Additionally, the entailment of a structural property enhances the representation and the discrimination powers of GNNs (Kipf and Welling, 2016; Velickovic et al., 2017).
@@
Early GNNs assume the network homophily, where nodes of similar features make connections based on social influence and/or selection theories. Spectral GNNs (Defferrard et al., 2016; Kipf and Welling, 2016) employ Laplacian smoothing called low-pass filtering to receive low-frequency signals from neighbor nodes. Consequently, these methods fail to adequately address heterophilous graphs (Zhu et al., 2020) and show dismal performance (Ma et al., 2021) in several heterophilous cases. To resolve this problem, numerous clever algorithms (spatial GNNs) have been proposed, including adjustment of edge coefficients (Velickovic et al., 2017; Brody et al., 2021) and aggregation of remote but highly similar nodes (Pei et al., 2020; Li et al., 2022). However, most of these methods assume positive or unsigned edges, which may impoverish the separation power (Yan et al., 2021). Recently, several studies (Chien et al., 2020; Bo et al., 2021) have applied message passing over negative (signed propagation) edges to preserve high-frequency signal exchanges. From the viewpoint of gradient flow, (Di Giovanni et al., 2022) shows that negative eigenvalues can intensify high-frequency signals during propagation. On the other hand, (Luo et al., 2021) prevents message diffusion by assigning zero-weights (blocking information) to heterophilic edges to alleviate local smoothing. Based on this observation, we raise a question: does signed propagation always attain the best performance on heterophilic graphs?
@@
To address the above question, we focus on the recent investigations (Ma et al., 2021; Yan et al., 2021) that provide valuable insights on message-passing. These studies mathematically analyze the effect of unsigned and signed messages on separability by comparing node feature distributions before and after message passing and aggregation. Furthermore, they reveal that signed messaging always enhances performance on heterophilic graphs. However, our empirical analysis (Fig. 3) shows that this assertion may not hold well under real-world graphs. This discrepancy arises because the prior work assumes binary class graphs (Aseem et al., 2023), impeding their extension to multi-class scenarios. In this paper, we deal with signed messaging on this condition. We first demonstrate that blind application of signed messaging to multi-class graphs incurs smoothing and performance degradation. To overcome the problem, we propose to utilize two types of calibration (Guo et al., 2017; Wang et al., 2021), which are simple yet effective in enhancing the performance of signed GNNs. In summary, our contributions can be described as follows:
@@
  • We hypothesize and prove that signed propagation reduces the discrimination power in certain cases while increasing prediction uncertainties. Expanding this, we propose novel strategies based on two types of calibration that are shown to be effective for both signed and unsigned GNNs.
@@
  • We perform extensive experiments with real-world benchmark datasets to validate our theorems. Our experiments demonstrate the effectiveness of the proposed techniques.
2. Related Work
@@
Graph Neural Networks (GNNs). Aggregating the information from adjacent nodes, GNNs have shown great potential under semi-supervised learning settings. Early study (Defferrard et al., 2016) focused on the spectral graph analysis (e.g., Laplacian decomposition) in a Fourier domain. However, it suffers from large computational costs as the graph scale increases. GCN (Kipf and Welling, 2016) reduced the overhead by harnessing the localized spectral convolution through the first-order approximation of a Chebyshev polynomial. Another notable approach is spatial-based GNNs (Velickovic et al., 2017; Brody et al., 2021; Morris et al., 2023; Barbero et al., 2023) which aggregate information in a Euclidean domain. Early spatial techniques have led to the development of many powerful schemes that encompass remote nodes as neighbors.
@@
GNNs on Heterophilic Graphs. Traditional message-passing GNNs fail to perform well in heterophilic graphs (Pei et al., 2020). To redeem this problem, recent studies have focused on handling disassortative edges (Choi et al., 2022; Zhao et al., 2023; Chanpuriya and Musco, 2022; Choi et al., 2023; Bi et al., 2024; Mao et al., 2024) by capturing node differences or by incorporating similar remote nodes as neighbors. More recently, H2GCN (Zhu et al., 2020) suggests the separation of ego and neighbors during aggregation. As another branch, selecting neighbors from non-adjacent nodes (Li et al., 2022), configuring path-level pattern (Sun et al., 2022), finding a compatibility matrix (Zhu et al., 2021), employing adaptive propagation (Wang et al., 2022), and choosing appropriate architectures (Zheng et al., 2023) have been recently proposed. Some methodologies change the sign of disassortative edges from positive to negative (Chien et al., 2020; Bo et al., 2021; Fang et al., 2022; Guo and Wei, 2022; Lee et al., 2023; Zheng et al., 2023; Yan et al., 2024; Li et al., 2024) while others assign zero-weights to disassortative edges (Luo et al., 2021). Even though (Baranwal et al., 2021) demonstrates that signed propagation can effectively handle binary class graphs, further investigation may be needed before applying the same techniques to multi-class graphs.
3. Preliminary
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$ be a graph with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges. The node attribute matrix is $X\in\mathbb{R}^{n\times F}$, where $F$ is the dimension of an input vector. Given $X$, the hidden representation of node features $H^{(l)}$ at the l-th layer is derived through message-passing. Here, node $i^{\prime}s$ feature is defined as $h_{i}^{(l)}$. The structural property of $\mathcal{G}$ is represented by its adjacency matrix $A\in\{0,1\}^{n\times n}$ and diagonal matrix $D$ of node degrees is derived from $A$ as $d_{ii}=\sum^{n}_{j=1}{A_{ij}}$. Each node has its label $Y\in\mathbb{R}^{n\times C}$ ($C$ represents the number of classes). Lastly, the global edge homophily ratio, $\mathcal{H}_{\mathcal{G}}$, is defined as:
(1) $\mathcal{H}_{\mathcal{G}}\equiv\frac{\sum_{(i,j)\in\mathcal{E}}1(Y_{i}=Y_{j})}{|\mathcal{E}|}$
Likewise, the local homophily ratio of node $i$ ($b_{i}$) is given by:
(2) $b_{i}\equiv\frac{{\sum^{n}_{j=1}A_{ij}\cdot 1(Y_{i}=Y_{j})}}{d_{ii}}$
@@ 4. Theoretical Analysis
We first discuss the mechanism of GNNs in terms of message-passing and parameter update (§4.1). Then, we introduce previous analyses of signed message passing on binary class graphs (§4.2) and highlight some misunderstandings revealed through an empirical study with real-world datasets (§4.3). Based on this observation, we extend the prior analysis to a multi-class scenario and point out some drawbacks of signed messaging in general environments with multiple classes (§4.4).
4.1. Message-Passing Neural Networks
@@
GNNs employ alternate steps of propagation and aggregation, during which the node features are updated iteratively. This is widely known as message-passing, which can be represented as follows:
(3) $\begin{gathered}H^{(l+1)}=\phi(\bar{H}^{(l+1)})\end{gathered}$
@@
Here, $H^{(0)}=X$ is the initial vector and $H^{(l)}$ is nodes’ hidden representations at the l-th layer. $\bar{H}^{(l+1)}=AH^{(l)}W^{(l)}$ is retrieved through message-passing ($A$) and we obtain $H^{(l+1)}$ after an activation function $\phi$. $W^{(l)}$ is trainable weight matrices shared by all nodes. The label prediction is produced by applying cross-entropy $\sigma(\cdot)$ (e.g., log-softmax) to $\bar{H}^{(L)}$, where the loss function is defined as below:
(4) $\mathcal{L}_{GNN}=\mathcal{L}_{nll}(Y,\widehat{Y})$
@@
The parameters are updated by computing negative log-likelihood loss $\mathcal{L}_{nll}$ between the predictions ($\widehat{Y}=\sigma(\bar{H}^{(L)})$) and true labels ($Y$). Since most GNNs assume homophily, they tend to preserve the low-frequency information (local smoothing) (Nt and Maehara, 2019). As a result, they fail to pinpoint the differences between heterophilic neighbors (Oono and Suzuki, 2019; Pei et al., 2020). To solve this, recent studies (Chien et al., 2020; Bo et al., 2021; Yan et al., 2021; Chen et al., 2022) focus on the extraction of high-frequency signals by flipping the sign of disassortative edges from positive to negative. In the following sections, we analyze the influence of information propagation from the perspectives of message-passing (Eq. 3) and parameter update (Eq. 4).
4.2. Using Signed Messages on Binary Classes
Before delving into the multi-class graphs, we assess the implications of three message-passing schemes (Eq. 7 - 9) under binary class graphs (Baranwal et al., 2021; Yan et al., 2021).
@@ 4.2. Using Signed Messages on Binary Classes
Message-passing (Signed propagation always has a better discrimination power than other propagation methods) For brevity, we take GCN (Kipf and Welling, 2016) as an example to inspect how message passing alters node feature vectors. Assuming a binary label ($y_{i}\in$0, 1}), let us inherit several assumptions and notations from (Yan et al., 2021) as follows: (1) For all nodes $i=\{1,...,n\}$, their degrees $\{d_{i}\}$ and features $\{h_{i}\}$ are i.i.d. random variables. (2) Every class has the same population. (3) With a slight abuse of notation, assume that $h^{(0)}=XW^{(0)}$ is the first layer projection of the initial node features. (4) The node feature of label $y_{i}$ follows the distribution ($\mu$ or $-\mu$) as,
(5) $\begin{gathered}\mathbb{E}(h^{(0)}_{i}|y_{i})\sim\begin{cases}\mu,&\text{if }y_{i}=0\\ -\mu,&\text{if }y_{i}=1\end{cases}\end{gathered}$
@@
We first introduce the equations from prior work (Yan et al., 2021) (Eq. 7 and Eq. 8), which demonstrate the updates in the expectations of an ego node $i$’s latent features $h^{(0)}_{i}$ after a single-hop application of message passing $h^{(1)}_{i}$ below:
(6) $h^{(1)}_{i}={h^{(0)}_{i}\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}{h^{(0)}_{j}\over\sqrt{(d_{i}+1)(d_{j}+1)}}$
(a) Binary class (C=2)
(b) Multiple classes (C=3)
Figure 1. We provide an example to illustrate the distribution of node features in (a) binary and (b) multi-class scenarios. Without loss of generality, in the case of having $C$ classes, the representation can be achieved by adding $C-2$ angles to the binary class. In both cases, one can easily infer that the scale of the aggregated neighbors is always smaller than $|\mu|$, considering the Laplacian- or attention-based aggregation methods (please see Appendix B-1 for more details)
@@
(Vanilla GCN) As illustrated in Figure 1a (binary class), we assume $h_{i}\sim N(\mu,0)$ for $y_{i}=0$. Otherwise, $h_{i}\sim N(-\mu,0)$. Let us assume that $d_{ij}=\sqrt{(d_{i}+1)(d_{j}+1)}$. Then, the updated node feature $\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})$ derived through the message-passing of vanilla GCN (Eq. 34) is as follows:
(7) $\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})=\left({1+(2b_{i}-1)d^{\prime}_{i}\over d_{i}+1}\right)\mu,\,\,\,d^{\prime}_{i}=\sum_{j\in\mathcal{N}_{i}}{1\over d_{ij}}$
@@
(Signed GCN) To define the mechanism of signed GCN, we need to assume an error ratio $e$ that stands for the accuracy ($1-e$) of configuring the sign of heterophilous edges. Considering that the sign of heterophilous nodes is flipped inaccurately with a probability ($e$), the update of the signed GCN can be defined as below:
(8) $\mathbb{E}_{s}(h^{(1)}_{i}|v_{i},d_{i})=\left({1+(1-2e)d^{\prime}_{i}\over d_{i}+1}\right)\mu$
@@
(Zero-weight GCN) Given an error ratio $e$, assigning zero weights to heterophilic edges leads to the following update in the expectation of feature distribution:
(9) $\mathbb{E}_{z}(h^{(1)}_{i}|v_{i},d_{i})=\left({1+(b_{i}-e)d^{\prime}_{i}\over d_{i}+1}\right)\mu.$
$\bullet$ Derivation of the above equations is in Appendix A-1 $\sim$ A-3.
For all message-passing schemes, if the coefficient of the right equations is smaller than 1, node feature vectors move towards the decision boundary and message-passing loses its discrimination power (Yan et al., 2021). Based on this, we compare the separability of signed GCN and zero-weight GCN below.
@@
Omitting the overlapping part of Eq. 8 and Eq. 9, their difference $Z=\mathbb{E}_{s}(\cdot)-\mathbb{E}_{z}(\cdot)$ can be derived using the error ratio and the local homophily ratio as,
(10) $\begin{gathered}Z=(1-2e)-(b_{i}-e)=1-e-b_{i},\end{gathered}$
@@
(a) $Z$ in binary class
(b) $Z$ in multiple classes
@@
Figure 2. We plot the Z in (a) Eq. 10 and (b) Eq. 18 to compare the discrimination powers of signed GCN and zero-weight GCNs. The red and blue colored parts indicate the regions where signed GCN and zero-weight GCN produce superior performances, respectively
Parameter update (Signed messages consistently improve separability in binary classification scenario) Let us assume that an ego node ($i$) and one of its neighbor nodes ($j$) are connected with a signed edge. Please ignore other neighbors and concentrate on the separation effect of signed propagation. To compute the loss function, we need to predict the label of the ego node ($i$) as follows:
@@
(11) $\begin{gathered}\widehat{Y}_{i}=\sigma(\bar{H}_{i}^{L})=\sigma\left(\frac{\bar{H}_{i}^{(L)}}{d_{i}+1}-\frac{\bar{H}_{j}^{(L)}}{\sqrt{(d_{i}+1)(d_{j}+1)}}\right)\end{gathered}$
@@
Assume that the true label of $i$ is $Y_{i}=k$. Then, the loss (Eq. 4) between $Y_{i}$ and the prediction ($\widehat{Y}_{i}$) can be computed as:
(12) $\mathcal{L}_{nll}(Y_{i},\widehat{Y}i)=-\log(\widehat{y}{i,k})$
@@
Accordingly, the update procedure for nodes $i$ and $j$ follow $\widehat{y}^{(t+1)}_{i,k}=\widehat{y}^{t}_{i,k}-\eta\nabla_{i}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})$ and $\widehat{y}^{(t+1)}_{j,k}=\widehat{y}^{t}_{j,k}-\eta\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})$, respectively. Here, $\eta$ is the learning rate and $\nabla$ represents a partial derivative. Then, the gradient of node $j$, $\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}$ follows:
@@
(13) $\nabla_{j}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}=-{1\over\widehat{y}_{i,k}}\cdot(\widehat{y}_{i,k}(1-\widehat{y}_{i,k})\underset{\nabla_{j}=\partial i/\partial j}{\underline{\textbf{(-1)}}})>0$
@@
Since the column-wise components of the projection matrices ($W$) used to compute $\bar{H}$ (Eq. 3) act as independent classifiers, as the training epoch ($t$) progresses, the probability of class $k$ for node $j$ ($\widehat{y}_{j,k}$) decreases compared to node $i$ since $\nabla_{i}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})<0.$
$\bullet$ More details of Eq. 13 can be found in Appendix A-4.